### v1

# Family Matters: Language Transfer and Merging for Adapting Small LLMs to Faroese

Abstract
We investigate how to adapt small, efficient LLMs to Faroese, a low-resource North Germanic language. Starting from English models, we continue pre-training on related Scandinavian languages, either individually or combined via merging, before fine-tuning on Faroese. We compare full fine-tuning with parameter-efficient tuning using LoRA, evaluating their impact on both linguistic accuracy and text comprehension. Due to the lack of existing Faroese evaluation data, we construct two new minimal-pair benchmarks from adapted and newly collected datasets and complement them with human evaluations by Faroese linguists. Our results demonstrate that transfer from related languages is crucial, though the optimal source language depends on the task: Icelandic enhances linguistic accuracy, whereas Danish boosts comprehension. Similarly, the choice between full fine-tuning and LoRA is task-dependent: LoRA improves linguistic acceptability and slightly increases human evaluation scores on the base model, while full fine-tuning yields stronger comprehension performance and better preserves model capabilities during downstream fine-tuning.
1 Introduction
While large language models (LLMs) excel in English and other high-resource languages, low-resource languages lag behind: model quality is tightly linked to data availability Robinson et al. (2023); Li et al. (2024), and even basic comprehension may fail Court and Elsner (2024). Coverage of these languages typically requires the largest available models, if it works at all.
In this paper, we focus on Faroese, a North Germanic language spoken by around 70,000 people, mostly in the Faroe Islands. Training data is scarce, with only 95 million words available in the deduplicated Fineweb-2 dataset Penedo et al. (2024). Faroese is particularly interesting for studying transfer from related languages. In the Nordic context, it has been described as the central Nordic language Torp (1998), reflecting its typological position in relation to the other Scandinavian languages. Derived from Old Norse, shaped by historical development and sociopolitical circumstances, modern Faroese shares characteristics with all the other Scandinavian languages.
We investigate methods for adapting small generative models to Faroese through continued pre-training, with particular emphasis on leveraging transfer from related languages. Prior work has demonstrated that syntactic similarity is a strong predictor of transfer success Chang et al. (2024), and for Faroese, encoder models benefit notably from transfer from Icelandic and other Scandinavian languages Snæbjarnarson et al. (2023). Beyond direct transfer, we explore parameter merging (Wortsman et al., 2022; Ilharco et al., 2023; Yadav et al., 2023), which enables the combination of models fine-tuned on different languages. This approach allows us to control the relative influence of each source language, a flexibility that would be more costly to achieve via multilingual training. Our goal is to balance the close linguistic similarity of Icelandic with the larger data resources available for mainland Scandinavian languages. To this end, we merge models trained on various Scandinavian languages and subsequently fine-tune the merged models on Faroese. We also compare two fine-tuning strategies for the continued pre-training: full fine-tuning and parameter-efficient fine-tuning with LoRA (Hu et al., 2022). Prior evidence suggests that full fine-tuning provides higher accuracy, while LoRA better preserves previously learned skills and diversity in text generation (Biderman et al., 2024). Based on this, we hypothesize that full fine-tuning will yield stronger linguistic performance in Faroese, but with a higher risk of losing general reasoning and knowledge abilities.
A major challenge for low-resource languages is the scarcity of evaluation data. To address this, we take two steps. First, we introduce two minimal-pair evaluation suites for small LLMs, combining pre-existing, newly collected, and adapted datasets: FoBLiMP, which targets syntactic and linguistic acceptability, and FoBCoMP, which targets text comprehension. Second, we perform expert human evaluations (carried out by Faroese linguists) on two tasks, one with and one without low-resource downstream task fine-tuning.
Our results indicate that transfer from related languages is crucial. Transfer from Icelandic yields the best performance in the linguistic probes in FoBLiMP, whereas higher-resource mainland Scandinavian languages contributes more to the text comprehension probes in FoBCoMP. Language merging shows promise for combining these advantages, but improvements are not consistent. Contrary to our expectations, LoRA outperforms full fine-tuning on linguistic acceptability in both automatic and human evaluations. In contrast, full fine-tuning achieves better results on comprehension tasks and coupled with downstream fine-tuning.
Our research questions are the following:
  1. RQ1
    What is the effect of transfer from (to various degrees) related languages?
  2. RQ2
    Does merging various related languages have a benefit over choosing the closest neighbor?
  3. RQ3
    What are the differences between fine-tuning all parameters of the base model and performing adaptation with LoRA?
2 Related Work
Language Adaptation
Post-hoc adaptation transfers knowledge from a high-resource or multilingual model to a new language, often using parameter-efficient methods. Early work focused on encoder models (Pfeiffer et al., 2020; Ansell et al., 2022; Ebrahimi and Kann, 2021). Yong et al. (2023) found that full fine-tuning works best for smaller models (e.g., 560M), while adapters are better for larger ones (up to 7.1B), but despite using a generative model, they do not evaluate on generative tasks. Razumovskaia et al. (2024) show that continued pre-training with LoRA (similar to our setup) improves linguistic quality of generations, though not few-shot ability. Instruction tuning on translated data is another option for language adaptation (Muennighoff et al., 2023b), but it is less reliable for small languages, and even correct translations can produce “translationese” (Gellerstam, 1986).
Selecting effective transfer languages is crucial for language adaptation. Chronopoulou et al. (2023) show that sharing adapter parameters across related languages improves machine translation for low-resource languages. Similarly, Faisal and Anastasopoulos (2022) find that linguistically informed adapter designs benefit unseen languages. In machine translation, transfer success often depends more on source corpus size and subword overlap than on broader linguistic similarity (Lin et al., 2019). Pairing a low-resource language with a typologically similar, higher-resource language can outperform using all available languages (Neubig and Hu, 2018). However, they focus on translation from a low-resource language; a setup that does not require generation in the small language.
Faroese Typology and Its Implications
Faroese is classified as Insular Scandinavian with Icelandic, while Norwegian, Swedish, and Danish are Mainland Scandinavian. This division reflects differences in lexicon, morphology, and syntax, though Faroese also shares traits with Mainland Scandinavian languages Thráinsson et al. (2012). Lexically, Faroese shares words and cognates with Icelandic and Danish, but no quantitative analysis exists Jacobsen (2021, 2022). Morphologically, it is more similar to Icelandic, with some overlap with Norwegian Torp (1998). Syntactically, Faroese is closer to Mainland Scandinavian languages Ussery and Petersen (2023); Debess (2017); Petersen (2010); Petersen and Heycock (2017); Sandøy (2005). Because of this mixed profile, identifying a single closest neighbour language is difficult. In our experiments, we fine-tune using data from all Scandinavian languages, with particular attention to Icelandic, to capture the different typological influences.
Evaluation for low-resource languages and small models
is challenging because there is little task-specific data. On the other hand, few-shot evaluations often focus on knowledge-intensive tasks or tasks that require structured outputs, making them out of scope for small models. Perplexity is the simplest measure of a model’s fit to a language, and only requires held-out text. In low-resource languages, however, fine-grained tokenization from common tokenizers reduces its usefulness (Oh and Schuler, 2024). An alternative is information parity, which compares the negative log-likelihood of a target-language text to its English version (Tsvetkov and Kipnis, 2024). This metric correlates with downstream performance but requires parallel data. Another option is translated benchmarks, but these miss culture-specific content (Chen et al., 2024), introduce translation errors, and produce “translationese”, making the target language artificially easier for the model.
Minimal pairs consist of two sentences that differ slightly, with one correct and one incorrect. Models are expected to assign higher probability to the correct sentence, capturing implicit preferences (Marvin and Linzen, 2018; He et al., 2025). Classical examples include subject-verb agreement (Linzen et al. (2016): “The key is on the table.” vs. “The key are on the table.”) and negative polarity items (Marvin and Linzen (2018): “No students have ever lived here.” vs. “Most students have ever lived here.”). BLiMP (Warstadt et al., 2020) provides a wide range of syntactic minimal pairs, with multilingual extensions like MultiBLiMP (Jumelet et al., 2025) including Faroese. COMPS (Misra et al., 2023) tests semantic knowledge (e.g., “A robin can fly.” vs. “A penguin/table can fly.”), and its multilingual extension (He et al., 2025) exists, though it does not cover Faroese.
Recent work for Faroese has focused on machine translation Scalvini and Debess (2024); Scalvini et al. (2025a); Simonsen and Einarsson (2024); Debess et al. (2025), combining automatic and human assessments. Automatic metrics such as BLEU or chrF prove shallow and fail to capture Faroese-specific nuances Scalvini et al. (2025b). Embedding-based metrics, such as BERTScore via FoBERT Snæbjarnarson et al. (2023), are emerging but limited and rarely validated against human judgment. Human evaluation remains essential. Beyond MT, small Faroese evaluation datasets exist for sentiment analysis Debess et al. (2024) and question answering Simonsen et al. (2025).
3 Experimental Setup
In this section, we introduce our experimental setup including model training (Section 3.1), merging (Section 3.2), collection of data for the automatic evaluation (Section 3.3), and the human evaluation process and data (Section 3.4). [^1] We make models and data available in a huggingface collection: huggingface.co/collections/jekunz/faroese-adaptation-68d6aea566b16d4f57180682.
3.1 Training
We use the two smaller SmolLM2 (Allal et al., 2025) models (135M and 360M parameters) as they are fully open, including their training data, and well-trained for their size on an English corpus. Although their size limits performance on knowledge-intensive tasks, it allows us to continually pre-train and compare them in different setups across substantial corpora. We experiment with two adaptation setups to answer RQ3: full-parameter fine-tuning and LoRA fine-tuning. We train for 5 epochs on the Faroese corpus (following Muennighoff et al. (2023a)’s scaling law for data-constraint training); we do not repeat data for other languages. Training details can be found in Table 6 in Appendix A. We train on the deduplicated Fineweb-2 (Penedo et al., 2024) portions for the Scandinavian languages, containing 27B tokens for Danish, 25B for Swedish, 30B for Norwegian-Bokmål, 1.6B for Icelandic, 495M for Norwegian-Nynorsk, and 95M for Faroese. Due to resource constraints, we limit the corpora for Swedish, Danish and Norwegian (Bokmål) to 4B tokens. We perform sequential continued pre-training for RQ1, first on an individual transfer language, then on the Faroese. We do not merge the data of the source languages with the Faroese data because this may introduce language mixing issues (Li et al., 2025).
3.2 Merging
Merging provides an efficient alternative to multi-task training by combining different fine-tuned models into one checkpoint, often improving generalization (Wortsman et al., 2022; Yadav et al., 2024). Task Arithmetic merging (Ilharco et al., 2023) computes task vectors as the difference between a fine-tuned model and its base, averages them, and adds the result back to the base model. TIES (Yadav et al., 2023) refines this by keeping only the most influential parameters, resolving sign conflicts, and using a disjoint mean to reduce interference. We apply TIES to languages instead of tasks, using Mergekit (Goddard et al., 2024): starting from SmolLM, we fine-tune on each language to obtain language vectors, merge them, and train the resulting model on Faroese. Since we continue pre-training on Faroese after merging and have limited resources, we select three promising merges that cover different language mixes and weightings: Mergeeq where we merge all five models equally (with pre-normalization weight 1 and density 0.5), Mergeis+ where we merge with bias towards Icelandic (weight 1 for Icelandic, 0.5 for all others, density 0.5), and Merge2, where we merge only two models: Icelandic and Danish (both with weight 1 and density 0.5), as Danish is the mainland Scandinavian language with the lowest perplexity after Faroese continued pre-training (see Table 3(b)).
3.3 Automatic Evaluation
We evaluate the perplexity on the validation set of the Faroese portion of Fineweb-2 both zero-shot and after continued pre-training on Faroese. In addition, we introduce two benchmarks: FoBLiMP for linguistic acceptability probes, and FoBCoMP for text comprehension probes. We report results on the original SmolLM models as a baseline.
FoBLiMP
To probe zero-shot linguistic skills, we use minimal pairs with one correct and one corrupted sentence, measuring the percentage of times the model assigns higher probability to the correct sentence. This collection is called FoBLiMP (Faroese Benchmark of Linguistic Minimal Pairs).
To evaluate subject-verb agreement, we use the Faroese portion of MultiBLiMP (Jumelet et al., 2025), containing 232 sentences. ScaLA (Nielsen, 2023) contains sentences corrupted by swapping or deleting words. Originally a binary classification task, we convert it to minimal pairs by realigning correct and incorrect sentences using Levenshtein distance ($\geq 0.85$), with unmatched samples added manually. Concatenating all subsets gives 552 pairs for flip_neighbours and 601 pairs for delete. GermDetect Michael and Horbach (2025) provide automatically corrupted sentences with verb placement errors. After removing pairs with no corruption, we obtain 2,026 pairs. As Faroese allows flexible word order, some corruptions are grammatical, but we conclude from an inspection that the original sentence is mostly more common. We also construct minimal pairs from a human evaluation in Scalvini et al. (2025a), where two raters annotated errors in English-to-Faroese translations from four models. We pair translations with an error difference of at least 2, keeping those with no more than four errors in the better translation and excluding translations containing foreign scripts. This yields 680 pairs.
FoBCoMP
Evaluating small LLMs in text comprehension is particularly challenging because evaluations often mix formal competence (e.g., grammar) with functional competence (e.g., following prompts) (Kydlíček et al., 2024). Limited fine-tuning data further complicates comparisons. To address this, we also use text comprehension probes in a minimal-pair format. We introduce a set of five probes, called FoBCoMP (Faroese Benchmark of Text Comprehension Minimal Pairs).
We adapt the Faroese news sentiment dataset (Debess et al., 2024) (original labels: positive, negative, neutral) into minimal pairs by adding a sentiment-bearing sentence (“Hetta er gott/ringt”). Neutral labels are excluded as initial experiments showed that words such as “neutral” are never the most probable choice. We evaluate sentence- and article-level samples, keeping only items annotators of the original dataset agreed on, resulting in 91 sentence-level (55 positive, 36 negative) and 84 article-level (51 positive, 33 negative) pairs. Using the same dataset, we filter GPT-4-assigned topic labels confirmed by a human. Minimal pairs consist of one correct topic and one incorrect topic (not assigned to the article), with related-topic pairs curated to make the task realistic (e.g., Local News vs. International News). This yields 234 topic classification pairs. We also adapt the extractive QA dataset FoQA (Simonsen et al., 2025) (2,000 Faroese question–context–answer triplets) into minimal pairs via two methods: (1) Dataset Shuffling: Replacing the correct answer passage with an incorrect but plausible passage from another sample within the context, matching token length. This creates 21,867 pairs. (2) GPT-4 Adversarial Answers: Generating one alternative incorrect answer per sample that is also a span in the dataset, matching token length when possible. Exact length matches occurred for 611 answers; deviations were an average of 1.69 tokens longer than the correct answers. This yields 2,000 pairs.
3.4 Human Evaluation
As we do not have evaluation sets to assess fine-grained properties for Faroese generation, we perform a human evaluation to assess output quality. The latter requires subjective judgment of linguistic quality, naturalness, and contextual appropriateness. As human evaluation is time-intensive, we focus on four 360M models, which perform among the best. These models vary along two dimensions: full tuning versus LoRA, and transfer from Icelandic versus Mergeis+. The evaluators are two native Faroese speakers, both trained linguists with extensive experience in NLP model evaluation.
Sentence Continuation
Given that we are working with small base models, we chose a simple generation task: sentence continuation. Models are prompted with sentences from a small manually compiled corpus derived from academic papers and local news (not in FineWeb2), with the last words removed and a trailing space. All models produce running text, enabling comparison of output quality. We evaluate outputs for linguistic quality across four subdimensions, scored 0–5: Lexical correctness (valid Faroese words and avoidance of lexical hallucinations), grammatical accuracy (morphological and syntactic correctness, including spelling and typography), Semantic coherence (meaningful, logically consistent content) and fluency/naturalness (native-like expression). The annotators each evaluated 400 continuations (100 prompts, 4 models) We apply a token cut-off of 100 and no penalty for incomplete last sentences.
Summarization
As zero-shot summarization did not result in reasonable summaries, we create a synthetic dataset for fine-tuning and manually evaluate the summaries generated by the models. The synthetic dataset is compiled with 150 authentic texts in the same domains as the evaluation set (academic, news, blog) paired with summaries generated by Claude Sonnet 4222For fine-tuning details, see Table 7 in Appendix A.. For evaluation, we selected 50 source texts (not part of training text) and generated summaries from all four models, resulting in 200 summary-source pairs. Two annotators conducted a blind evaluation using two criteria: task completion (to what extent did the model solve the task) and linguistic quality, each on a 0–5 scale. We apply a token cut-off of 400 and no penalty for incomplete final sentences.
4 Results and Discussion
4.1 Benchmarks
We first give an overview of the benchmarks results with respect to their difficulty and reliability. FoBLiMP results are shown in Table 1. Models perform well on most linguistic acceptability probes, suggesting these tasks are relatively easy. The main exception is Translation Pairs, where scores are lower, likely due to noise as translation error counts do not always reflect linguistic quality: an error can reflect unrelated aspects such as incorrect content compared to the source sentence.
Full LoRA 135M 360M 135M 360M En 95.25 96.55 96.12 96.55 +Da 96.98 95.25 98.27 96.12 +Is 97.41 99.13 97.41 98.27 +NoB 96.12 100.0 96.98 96.98 +NoN 96.12 98.27 95.68 97.41 +Sv 96.12 98.27 96.98 96.98 Mergeeq 97.41 98.27 96.55 95.25 Mergeis+ 97.84 97.84 97.41 96.55 Merge2 97.84 97.41 97.41 99.13
(a) Subject-verb agreement (MultiBLiMP).
Baseline: 66.81 (135M), 70.68 (360M)
Full LoRA 135M 360M 135M 360M 93.65 94.56 92.75 93.84 94.74 94.02 95.10 96.01 93.84 95.65 94.38 96.01 94.38 95.10 95.28 95.10 93.65 94.92 94.02 94.20 95.28 94.20 95.65 94.92 94.56 94.38 96.73 95.47 94.20 95.28 95.10 95.10 95.28 95.10 93.84 95.28
(b) ScaLA: flip_neighbors.
BL: 59.96 (135M), 62.50 (360M)
Full LoRA 135M 360M 135M 360M 87.18 93.17 84.52 88.51 90.84 93.01 90.18 92.84 89.51 94.00 90.18 95.34 90.34 92.34 90.68 94.50 88.51 93.17 89.01 91.01 89.85 93.34 90.51 93.17 90.18 94.34 89.35 93.01 92.01 94.00 89.01 93.17 90.68 93.51 90.34 91.84
(c) ScaLA: delete.
BL: 65.39 (135M), 67.38 (360M)
Full LoRA 135M 360M 135M 360M En 93.48 95.36 92.54 94.66 +Da 94.27 95.11 94.91 95.75 +Is 94.27 95.16 94.57 95.75 +NoB 95.06 95.31 94.91 95.55 +NoN 93.97 94.91 94.47 95.01 +Sv 94.61 95.85 95.36 95.80 Mergeeq 94.66 95.80 95.16 95.31 Mergeis+ 94.57 95.60 95.06 95.26 Merge2 94.52 95.85 94.37 95.60
(d) Verb placement (GermDetect).
Baseline: 52.22 (135M), 57.94 (360M)
Full LoRA 135M 360M 135M 360M 70.44 75.00 68.97 74.70 75.58 78.23 75.88 76.76 73.23 77.35 75.44 78.67 75.14 76.17 75.44 77.79 70.73 75.44 71.17 74.70 76.32 76.17 75.88 74.11 76.02 75.88 75.00 75.44 76.91 77.20 74.55 73.97 74.55 77.20 75.44 75.44
(e) Translation Pairs.
BL: 42.64 (135M), 46.32 (360M)
Full LoRA 135M 360M 135M 360M 88.00 90.92 86.78 89.65 90.48 91.12 90.86 91.49 89.65 92.25 90.39 92.80 90.20 91.78 90.65 91.98 88.59 91.34 88.87 90.46 90.43 91.55 90.87 90.99 90.56 91.73 90.55 90.89 91.10 91.98 90.22 90.81 90.57 91.81 90.28 91.45
(f) Mean of the individual FoBLiMP scores in 1(a)-1(e).
Table 1: Linguistic probes on datasets included in FoBLiMP: Percentage of samples where a higher probability was assigned to the original than to the corrupted sample.
FoBCoMP results are shown in Table 2. Scores are lower and more mixed than for FoBLiMP, reflecting the tasks’ difficulty for small models. In sentiment analysis, 135M LoRA models perform poorly, with little improvement over the base model, and full fine-tuning shows similar limitations at the article level, indicating 135M models are too small for zero-shot text comprehension. Topic classification results are challenging to interpret due to small, variable data; individual dataset results should generally be interpreted cautiously. For extractive QA, results vary based on the setup. On the shuffled dataset, transfer offers limited gains. On the harder GPT-4–picked answers however, transfer improves scores, especially for 360M models.
Full LoRA 135M 360M 135M 360M En 60.43 68.13 60.43 72.52 +Da 72.52 75.82 61.53 76.92 +Is 63.73 75.82 60.43 69.23 +NoB 75.82 71.42 60.43 72.52 +NoN 76.92 74.72 67.03 71.42 +Sv 70.32 68.13 60.43 69.23 Mergeeq 75.82 72.52 60.43 62.63 Mergeis+ 68.13 80.21 60.43 63.73 Merge2 76.92 75.82 62.63 65.93
(a) Binary sentiment analysis (Sentences).
Baseline: 60.43 (135M), 60.43 ( 360M)
Full LoRA 135M 360M 135M 360M 60.71 70.23 60.71 73.80 61.90 71.42 60.71 70.23 61.90 71.42 60.71 71.42 63.09 69.04 60.71 70.23 65.47 63.09 64.28 69.04 70.23 65.47 60.71 63.09 61.90 63.09 60.71 60.71 64.28 69.04 60.71 63.09 69.04 73.80 63.09 61.90
(b) Bin. sentiment (Articles).
BL: 59.52 (135M), 60.71 ( 360M)
Full LoRA 135M 360M 135M 360M 77.35 84.18 67.94 67.94 75.21 82.47 71.79 70.94 79.05 78.20 71.79 76.06 68.80 80.76 72.22 74.78 60.68 79.05 69.65 61.11 63.67 73.50 65.38 80.34 66.23 77.77 68.37 70.94 75.64 82.90 59.82 73.93 75.21 78.63 57.26 72.64
(c) Topic Classification (Articles).
BL: 54.70 (135M), 60.25 ( 360M)
Full LoRA 135M 360M 135M 360M En 75.35 84.37 69.57 86.40 +Da 72.86 86.74 69.13 82.93 +Is 75.64 85.46 71.80 86.87 +NoB 72.04 84.16 70.00 86.60 +NoN 72.89 85.05 74.41 86.12 +Sv 70.82 85.12 70.42 86.10 Mergeeq 63.47 83.87 67.43 78.08 Mergeis+ 70.06 83.33 70.63 85.77 Merge2 72.25 85.14 70.19 82.43
(d) Extractive QA (Shuffled DS).
Baseline: 63.05 (135M), 68.94 (360M).
Full LoRA 135M 360M 135M 360M 54.40 55.25 52.40 54.85 57.65 66.70 59.70 66.35 49.20 65.05 55.35 67.35 55.85 59.15 54.35 64.60 54.65 62.60 54.30 60.45 56.20 63.45 53.00 65.25 53.20 63.70 50.85 60.05 56.60 62.05 51.85 62.70 56.35 67.65 58.75 66.00
(e) Extractive QA (LLM gen.).
BL: 32.90 (135M), 35.40 (360M).
Full LoRA 135M 360M 135M 360M 65.64 72.43 62.21 71.10 68.02 76.63 64.57 73.47 65.89 75.19 64.01 74.18 67.12 72.90 63.54 73.74 66.12 72.90 65.93 69.62 66.24 71.13 61.98 72.80 64.12 72.19 61.55 66.48 66.94 75.50 60.68 69.84 69.95 76.21 62.38 69.78
(f) Mean of the individual FoBCoMP scores in 2(a)-2(e).
Table 2: Text Comprehension Probes on datasets included in the FoBCoMP benchmark; Debess et al. (2024) and FoQA. Percentage of samples where a higher probability was assigned to the correct than to the incorrect sample.
Full LoRA 135M 360M 135M 360M En 73.61 58.27 73.61 58.27 +Da 50.98 40.45 63.23 44.38 +Is 38.77 30.09 40.05 30.54 +NoB 48.59 44.81 56.25 44.51 +NoN 61.58 50.28 68.02 56.62 +Sv 69.96 57.02 78.22 61.43 Mergeeq 103.45 42.40 182.21 338.30 Mergeis+ 94.95 40.16 68.81 62.88 Merge2 65.28 40.39 78.46 146.20
(a) Before continuing training on Faroese (zero-shot).
Full LoRA 135M 360M 135M 360M 4.98 3.75 5.51 4.48 4.19 3.56 4.25 3.55 4.44 3.48 4.53 3.53 4.22 3.63 4.26 3.56 4.60 3.66 4.90 4.08 4.26 3.60 4.21 3.58 4.08 3.41 4.61 3.93 4.08 3.41 4.58 3.77 4.22 3.49 4.56 3.80 arX
(b) After continuing training on Faroese.
Table 3: Average per-token perplexity on the Fineweb-2 evaluation set.
4.2 RQ1: Transfer Languages are Important
Across all benchmarks, initializing Faroese models with a Scandinavian transfer language improves performance compared to English-only models. The choice of transfer language, however, matters.
Perplexity
Models adapted via a Scandinavian transfer language consistently show lower perplexities than those adapted directly from English. The difference is larger for smaller models and for LoRA models, suggesting that LoRA benefits more from better-initialized parameters, consistent with Biderman et al. (2024) who find full tuning more sample-efficient than LoRA in domain adaptation.
Table 3(a) shows zero-shot perplexities before Faroese adaptation. Icelandic performs best, followed by Danish and Norwegian-Bokmål; Swedish is barely better than English-only. This likely reflects script overlap: Faroese shares most letters with Icelandic, some diacritics with Danish and Norwegian, but none beyond the English alphabet with Swedish. After Faroese adaptation (Table 3(b)), Icelandic remains best for larger models, while for smaller models, higher-resource mainland Scandinavian languages outperform Icelandic (except for Norwegian-Nynorsk, the lowest-resorce language).
FoBLiMP
While Icelandic does not always yield the best language modeling performance (perplexity after Faroese tuning), it achieves the highest scores on MultiBlimp (subject-verb agreement), as expected since Icelandic has subject-verb agreement, unlike the mainland Scandinavian languages.
Across all linguistic probes, results are mixed. Icelandic shows a small advantage (Figure 3), but it is not the top language for most task-model pairs. Mean scores across FoBLiMP tasks (Table 1(f)) show Icelandic performs best in 2 of 4 setups, making it the best individual language but not an undisputed leader. In particular, models adapted first to Icelandic do not consistently outperform those adapted to Danish or Norwegian Bokmål. Nynorsk performs worst, with no wins (Figure 3) and the lowest mean scores (Table 1(f)), likely due to limited data. This suggests that other features can compensate for lower surface similarity: the higher-resource mainland Scandinavian generally match Icelandic (except in SVA).
FoBCoMP
While Icelandic remains a strong transfer language, Nynorsk has the same win count (4), despite poor results in linguistic probes. Three of Nynorsk’s wins are from the LoRA 135M model, while other Nynorsk models perform worse, as reflected in the mean scores over FoBCoMP tasks (Table 2(f)). Table 2(f) shows Danish surpassing Icelandic in FoBCoMP: Danish wins against Icelandic in 3 of 4 aggregated cases, Nynorsk in 2/4, Bokmål and Swedish in 1/4, and English-only never. We conclude that results are very mixed.
In general,
Icelandic often provides the strongest transfer. It gives the best perplexity before Faroese adaptation and clear advantages in FoBLiMP probes like subject-verb agreement, supporting our choice of Icelandic for human evaluation. Figure 1 shows that choosing Icelandic over English gives almost consistent improvements for FoBLiMP and in many cases gains for FoBCoMP.
Figure 1: Win rates of Icelandic over English only. FoBLiMP (left), FoBCoMP (right).
Figure 2 illustrates how often all transfer setups outperform English, indicating that any additional transfer language is better than none. For FoBLiMP, especially in 360M models, transfer is crucial, while for FoBCoMP, the best language choice is less clear.
Figure 2: Win rates of all transfer setups over English only (i.e., any Scandinavian language is better than none. FoBLiMP (left), FoBCoMP (right).
Comparing single languages for the transfer (Figure 3), the choice is less clear: Icelandic has 14 wins and Danish 13. For FoBLiMP, Icelandic leads with 9 wins versus 7, and 4 ties; for FoBCoMP, Danish leads with 9 wins versus 6, and 5 ties.
Figure 3: Win rates of transfer languages across models and adaptation setups. In case of a tie, we count both. FoBLiMP (left), FoBCoMP (right).
4.3 RQ2: Effects of Merging
Perplexities
We see in Table 3(b) that merging all transfer languages (setups Mergeeq and Mergeis+) leads to the lowest perplexities in full-parameter fine-tuning, suggesting that with sufficient learning capacity — and at least for language modelling — models can benefit from a mixture of related languages. For LoRA, however, the opposite holds: the fewer languages merged, the better the results, with the setups without merging performing best. Full fine-tuning appears more robust to language merging. Merging LoRA-fine-tuned models may generally be more problematic than merging fully fine-tuned models, as other work has also shown, possibly due to a weaker degree of representation alignment (Stoica et al., 2024). Zero-shot perplexities (Table 3(a)) in particular for Mergeeq are exploded; the more is merged the higher the perplexities. But this, interestingly, still results in a better initialization for full fine-tuning.
FoBLiMP
In Table 1(f), we see that merges are beneficial for the full tuning setup: For the 135M model, the best 3 models are the 3 merges. For the 360M model, Icelandic is best, but followed by 2 merges. For LoRA however, the situation is different: For the 135M model, merges are within the same range as models with individual languages, while for the 360M models, the scores of merges are lower than for 3 out of 5 single languages.
FoBCoMP
The Merge2 setup, which combines Icelandic and Danish equally, has four wins in Table 2(f) and is clearly the best of the merges. Full fine-tuning benefits particularly the Merge2 setup, which is best overall setup for 135M and runner-up for 360M. For LoRA, single-language setups however outperform merges. It is better than Icelandic only for both full fine-tuning models, while Icelandic only is better for both LoRA models, again demonstrating that merges show promise for full fine-tuning but are less suitable for LoRAs.
Can Merging Improve Specific Aspects?
In the preceding paragraphs, we found that merges are sometimes competitive but most often outperformed by individual languages: We found that Icelandic tends to give stronger gains on the linguistic probes in FoBLiMP, while Danish seems more helpful for comprehension probes in FoBCoMP. This raises the question of whether the two languages can complement each other, with Icelandic adding linguistic strengths and Danish adding text comprehension skills. We therefore ask: Do merges add linguistic capabilities to transfer from Danish only? To test this, we compare Danish-only with merges on FoBLiMP (Table 1). Comparing each merge pairwise against Danish within otherwise equal setups, the merge wins in 30/60 cases and Danish in 29 cases, and in one case, there is a tie. To see the potential of merging, we also compare the best merge against Danish. The best merge outperforms Danish in 14/20 cases, while Danish performs better in 6 cases. These results suggest that merging can indeed add linguistic capabilities to Danish models. Exploring this further is a promising direction for future work.
Finally, we ask: Do merges add comprehension capabilities to Icelandic models? For FoBCoMP, we make a pairwise comparison of Icelandic with the merged models (based on Table 2). Overall, the merges do not perform better: Icelandic wins in 38 out of 60 comparisons, while the merges win 16 times, with 6 ties. Focusing only on the best merge in each case, the results are evenly split, with 10 wins for the merge and 10 wins for Icelandic. This suggests that merges can sometimes match or improve on Icelandic only, but the potential is smaller than for linguistic capabilities.
Human Ratings
in the sentence continuation task, (Table 5) show that the models trained on Icelandic only and the Mergeis+ model are very close: For LoRA, Mergeis+ achieves the highest overall score (3.436 versus 3.396), while for full tuning, the Icelandic model achieves higher average scores (3.413 versus 3.361). Icelandic LoRA performs best on the lexical level but is weaker in grammar, semantics and fluency. Interestingly, for semantics, the merged models score slightly higher in both cases, which is partially in line with results previously discussed in this section, where the Icelandic models were comparatively weak on comprehension tasks. This indicates that that while lexical knowledge can be effectively acquired through pre-training on Icelandic, other skills can benefit from the broader exposure provided by the merging approach. In the summarization task however (Table 5), scores of the Icelandic model are higher in both cases and across both linguistic quality and task completion, which could indicate the opposite.
Model Overall Lexical Grammar Semantics Fluency LoRA-Mergeis+ $3.436$ $3.865$ $3.645$ $3.015$ $3.216$ LoRA-Is+ $3.396$ $3.983$ $3.587$ $2.891$ $3.126$ Full-Is+ $3.413$ $3.940$ $3.600$ $2.955$ $3.160$ Full-Mergeis+ $3.361$ $3.924$ $3.535$ $2.934$ $3.051$
Table 4: Sentence Continuation scores (averages over both annotators). Scale 0–5, higher is better. Pearson’s correlation coefficient between annotators: $r=0.546$.
Model Task Completion Linguistic Quality Full-Is+ $3.01$ $4.11$ Full-Mergeis+ $2.92$ $3.82$ LoRA-Is+ $1.20$ $2.02$ LoRA-Mergeis+ $1.03$ $1.78$
Table 5: Summarization scores (average over both annotators). Scale 0–5, higher is better. Pearson correlation coefficient between annotators: $r=0.879$.
4.4 RQ3: Full Fine-Tuning versus LoRA
Overall, full fine-tuning yields higher results than LoRA across most evaluations, but particularly in merging setups (as discussed in 4.3). This advantage is especially visible in the text comprehension tasks. LoRA shows its relative strength in linguistic acceptability transfer, especially for 135M models.
Perplexity
Full fine-tuning consistently outperforms LoRA in reducing perplexity, showing that the increased learning capacity is crucial for core language modeling. The effect is even stronger when multiple transfer languages are merged, where LoRA consistently underperforms.
FoBLiMP
In Table 1(f), we see that for the smaller 135M model, LoRA outperforms full fine-tuning in 5 out of 6 cases for single languages, while for the 360M model the results are balanced (3 wins each). Looking at all individual tasks in Figure 4, LoRA again has a clear advantage for the 135M model (34 wins vs. 23 for full fine-tuning, with 3 ties), but for the 360M model both methods perform similarly, with full tuning having a slight edge. For merges, however, both the mean scores in Table 1(f) and the win rates in Figure 4 indicate that full fine-tuning is better overall. These findings highlight that LoRA performs surprisingly well for the 135M model, capturing syntactic transfer efficiently despite being more parameter-efficient. Interestingly, this contradicts our initial assumption that higher learning capacity is more important for acquiring linguistic skills, while preventing catastrophic forgetting would primarily benefit comprehension tasks.
Figure 4: Wins of LoRA vs. full FT for model twins on FoBLiMP. No merges (left), merges (right).
FoBCoMP
results for full tuning and LoRA are compared in Figure 5. Full tuning consistently outperforms LoRA: full tuning achieves 35 wins compared to 22 for LoRA (with 3 ties; see Figure 5). This may seem counterintuitive, as fine-grained language skills are not obviously critical for comprehension tasks. However, it is possible that full tuning provides a more stable adaptation process, which benefits performance on these datasets.
Figure 5: Wins of LoRA vs. full FT for model twins in FoBCoMP. No merges (left), merges (right).
Model size effects
Differences between full tuning and LoRA are mainly driven by the 135M model. For the 360M model, margins between both approaches are small, suggesting that larger models are more robust to adaptation method choice.
Human Ratings
for sentence continuation (Table 5) give mixed results: For merges, LoRA models score higher than fully tuned models (3.436 versus 3.361), while for Icelandic-only models, the fully tuned model scores higher (3.413 versus 3.396). However, for summarization (Table 5), the results are very clear: Fully tuned models score much higher. The fully tuned Icelandic model achieves the highest performance of all with a task completion score of 3.01 and linguistic quality score of 4.11, while LoRA models perform substantially worse, with LoRA-Mergeis+ scoring only 1.03 in task completion and 1.64 in linguistic quality. This strong difference suggests that full tuning provides a better, or at least more stable, surface for preserving linguistic skills during downstream task fine-tuning. Interpretations that full tuning benefits the higher-level organizational task of summarization may be possible but should be done cautiously, given the noise added by the task fine-tuning setup with very little data that we use for summarization.
4.5 Qualitative Observations
The low semantic scores across models highlight a limitation in adaptation for low-resource languages. While the models acquire competency in surface-level linguistics — producing valid Faroese vocabulary, maintaining grammatical structures, and achieving natural-sounding fluency — they struggle significantly with generating meaningful, coherent content. This reflects a fundamental challenge in language model training: while syntactic and lexical patterns can be learned from limited data through transfer from related languages, semantic understanding — which requires world knowledge — is a challenge, particularly in small models. The example in Figure 6 highlights this: although the models produce text that appears fluent, they often fail to convey coherent ideas, limiting their utility for applications requiring content generation.
Many summarization outputs exhibit mixes between languages, as in the example in Figure 7. This was not the case for sentence continuation outputs, which indicates that the low-resource tuning destroyed some of the models’ linguistic abilities.
Faroese: Translation:
Input: Børn millum fimm og 11 ár verða nú koppsett fyri koronu nú Input: Children between five and 11 years are now being vaccinated for coronavirus now
Continuation: á døgninum. Síðani frísparksmeginreglurnar vórðu settar í gildi 24. apríl hevur eingin smitta verið staðfest, síðan stigtakarin til spæliskeiið - Danmark - varð settur at hava tamarhald á korona-smittuni týsdag Continuation: during the day. Since the free kick regulations were put into effect on April 24th, no infection has been confirmed, since the initiator of the play workshop - Denmark - was set to have control over the corona infection on Tuesday
Figure 6: Example: Semantic nonsense in outputs.
Example of summary output:
Samandráttur er ein lokaliserendre i fjórðhálsparafjöllum fyrir fjölkvangna forfælja í Føroyskaflokkum til víkjandi barna uppaling er lutfalsliga sterk í Føroyum.
Figure 7: Language mixing in outputs. The model starts in Faroese, then uses language similar to Danish or Norwegian (in italics), then language resembling Icelandic (underlined), then, again, Faroese.
4.6 Limitations and Future Work
The small amount of evaluation data for Faroese is a major limitation, and even the human study covers relatively few samples. Summarization results in particular should be viewed with caution as the small validation set made hyperparameter tuning difficult, and in particular the LoRA models may not have been trained optimally.
Valuable directions for future work are to investigate how these findings change with model sizes and after instruction tuning, and if adaptive or data-driven merging strategies can lead to better transfer from multiple languages. Multilingual studies could reveal whether the same patterns hold for other low-resource languages in other families.
5 Conclusion
We studied how to adapt small LLMs to Faroese by transferring from related Scandinavian languages, merging models trained on these languages, and comparing full and parameter-efficient fine-tuning. We found that transfer from related languages is essential, but the best source varies. Icelandic gave the strongest gains in linguistic accuracy, while Danish was more helpful for comprehension. This suggests that practical applications should draw on multiple sources rather than relying only on the closest relative. Merging showed potential, although the benefits were inconsistent and more sensitive under LoRA than full fine-tuning. This indicates that merging can add complementary strengths, but requires careful design. LoRA proved effective for improving linguistic acceptability, while full fine-tuning performed better on comprehension-heavy tasks. Full fine-tuning also provided a stronger base for downstream fine-tuning on summarization in low-resource settings. Overall, our results show that transfer from related languages is key for adapting LLMs to low-resource settings, and that the optimal tuning strategy depends on priorities and target application.
Acknowledgments
This research was supported by TrustLLM funded by Horizon Europe GA 101135671. The computations were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre and by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Appendix A Training Details
Parameter Value
Optimizer AdamW (Loshchilov and Hutter, 2019)
Scheduler Cosine with 5% warmup
Batch size 256 (effective)
Context window 8192 tokens
Learning rate Full fine-tuning: $5\times 10^{-4}$
LoRA: $8\times 10^{-4}$
(also tested: $5\times 10^{-5}$$1\times 10^{-3}$)
Hardware 1 node, 4 or 8 A100 40GB GPUs
LoRA rank 256
LoRA $\alpha$ 512
LoRA #parameters 57.5M (135M); 102M (360M)
Training epochs 5 (Faroese corpus), 1 (all other languages)
Total compute 5,000 A100 (40GB) hours
Table 6: Training hyperparameters and setup for continued pre-training.
Parameter Value
Optimizer AdamW (Loshchilov and Hutter, 2019)
Scheduler Cosine with 0.1 warmup
(also tested without scheduler)
Batch size 8
Context window 8192 tokens
Learning rate $5\times 10^{-5}$
(also tested: $5\times 10^{-4}$$1\times 10^{-3}$)
Dropout 0.1
(also tested: 0)
Hardware 1 A100 40GB GPU
LoRA rank 16
(also tested: 8)
LoRA $\alpha$ 32
(also tested: 16)
Training epochs 50 (for the full dataset, 151 samples)
Tuning split 135 training / 16 validation samples
Prompting Faroese, minimalistic setup indicating start of text and summary
Table 7: Hyperparameters and setup for summarization fine-tuning experiments in the human evaluation.